# On Conjecture of Binomial Edge Ideals of Linear Type

### v1

@@ Abstract.
An ideal $I$ of a commutative ring $R$ is said to be of linear type when its Rees algebra and symmetric algebra exhibit isomorphism. In this paper, we investigate the conjecture put forth by Jayanthan, Kumar, and Sarkar in [14] that if $G$ is a tree or a unicyclic graph, then the binomial edge ideal of $G$ is of linear type. Our investigation validates this conjecture for trees. However, our study reveals that not all unicyclic graphs adhere to this conjecture. Furthermore, we provide an explicit description of the defining ideal of the Rees algebra of binomial edge ideals of trees and demonstrate that these algebras are Cohen-Macaulay. Finally, we conclude the equality between symbolic powers and ordinary powers of binomial edge ideals of trees.
@@ Introduction
is called binomial edge ideal of the graph $G$. We call the binomial $f_{ij}$ associated to an edge $\{i,j\}\in E(G)$ as an edge binomial. In [10], Herzog et al. observed that the binomial edge ideal appears naturally in the study of conditional independence ideals, providing a valuable tool to investigate robustness theory in the context of algebraic statistics. In recent years, researchers have investigated various algebraic properties, including radicality, primality, Gröbner bases, primary decomposition, Cohen-Macaulayness, symbolic powers, among others, of binomial edge ideals in relation to their underlying combinatorial properties; see for partial list [6, 7, 10, 18, 19, 23, 24] and the references therein.
@@
[14, Conjecture 4.17. (a)] If $G$ is a tree or a unicyclic graph, then $J_{G}$ is of linear type.
@@
Secondly, when $G$ is a tree, we study symbolic and ordinary powers of $J_{G}$, and the Cohen-Macaulay property of $\mathcal{R}(J_{G})$. The study of symbolic powers is one of central topics in commutative algebra. In general symbolic powers and ordinary powers of an ideal do not coincide. There are some cases where it coincides. For the case of binomial edge ideals, Ohtani obtain the equality for complete multipartite graphs and Ene and Herzog obtain the equality for closed graphs (cf. [6, 20]). In [7], the authors showed that $\mathcal{R}(J_{G})$ is Cohen-Macaulay when $G$ is a closed graph. In this paper, we obtain the equality for trees by showing initial ideal of binomial edge ideals of trees are normally torsionfree, see Theorem 5.1. An ideal is said to be normally torsionfree if all its powers has same associated primes (cf. [25, Definition 5.2]). Furthermore, we shown that the Rees algebra binomial edge ideals of trees are Cohen-Macaulay.
This paper is organized as follows. The following section recalls definitions, notations, and known results required in the rest of the paper. In Section 2, we introduce and study $p$-sequence ideals. In Section 3, we shown that binomial edge ideals of trees and class of unicyclic graphs form a $p$-sequence. We study the Rees algebra of binomial edge ideals of trees in Sections 4 and 5. In Section 6, we present a counterexample to Conjecture A for a class of unicyclic graphs.
@@ 1. Preliminaries
  • [18, Definition 3.1] Let $e^{\prime}=\{i,j\}\notin E(G)$ be an edge in $G\cup\{e^{\prime}\}$. Then $G_{(e^{\prime})}$ is the graph on vertex set $V(G)$ and edge set
@@
[14, Lemma 4.1] [10, Corollary 2.2] Let $G$ be a simple graph and $f$ be an element of $S$. Then $J_{G}:f=J_{G}:f^{n}$ for all $n\geq 2$.
@@
[18, Theorem 3.4] Let $G$ be a simple graph. Let $e=\{i,j\}\notin E(G)$ be a bridge in $G\cup\{e\}$. Then $J_{G}:f_{e}=J_{G_{e}}$.
@@
[18, Theorem 3.7] Let $G$ be a simple graph and $e=\{i,j\}\notin E(G)$ be an edge. Then
@@
[24, Corollary 2.5] For any bipartite graph $G$, one has $\beta_{i,i+1}(S/J_{G})=0$, for all $i$.
@@
[13, Definition 1.1] Let $A$ be a commutative ring. Set $z_{0}=0$. A sequence of elements $z_{1},\ldots,z_{m}$ in $A$ is said to be a $d$-sequence if it satisfies the conditions:
@@
[14, Proposition 4.10] Let $A$ be a polynomial ring over a field $\mathbb{K}$. Let $I\subset A$ be a homogeneous almost complete intersection ideal. Then $I$ is generated by a homogeneous $d$-sequence $z_{1},\ldots,z_{h+1}$ such that $z_{1},\ldots,z_{h}$ is a regular sequence, $h=\operatorname{ht}(I)$.
@@
[22, Definition 1.1] A relation on an ordered sequence of elements $z_{1},\ldots,z_{n}$ of a commutative ring $A$ is a form of positive degree in the graded polynomial ring $R=A[X_{i}\mid 1\leq i\leq n]$, where elements of $A$ have degree 00 and each $X_{i}$ has degree $1$, that yields 00 when evaluated at $z_{1},\ldots,z_{n}$. Let $\Lambda_{j}$ denote the ideal of $R$ generated by the relations on $z_{1},\ldots,z_{n}$ of degree at most $j$. Let $r$ be the least positive integer such that $\Lambda_{r}=\cup_{j\geq i}\Lambda_{j}$. The invariant $r$ is called the relation type of the ideal $I=(z_{1},\ldots,z_{n})$. An ideal of relation type $1$ is said to be of linear type.
@@
[5, p. 257] Let $A$ be a commutative ring. An ordered sequence of elements $z_{1},\ldots,z_{n}$ of a ring $A$ is said to be a sequence of linear type if each of the ideal $(z_{1},\ldots,z_{i})$ is of linear type for all $1\leq i\leq n$.
@@
    Consider a sequence of elements $\{x_{1}x_{3}x_{4}x_{5},x_{1}^{2}x_{2}x_{6},x_{1}^{2}x_{2}^{2}x_{3}x_{5}\}$ in $\mathbb{K}[x_{1},\ldots,x_{6}]$. In the specified order, this sequence readily forms a $d$-sequence. However, no permutation of this sequence constitutes a $p$-sequence. This assertion is verified by Computation 6.1.
@@
Suppose the sequence of elements $z_{1},\ldots,z_{n}$ in $S$ form a $d$-sequence. Then, it satisfies Equation (2.3), see [12, Proposition 3.1]. However, this property does not hold for $p$-sequences. For instance, consider a graph $G$ with edge set
@@ 4.1. Defining ideal of Rees algebra
The subsequent result follows from Theorem 3.1 and [14, Theorem 3.2], where we obtain a minimal generating set for the defining ideal of $\mathcal{R}(J_{G})$ when $G$ is a tree.
@@
5. Cohen-Macaulayness
In this section, we prove that the initial ideal of binomial edge ideals of trees are normally torsionfree. Additionally, we show that the symbolic powers and ordinary powers of binomial edge ideals of trees coincide. Finally, we examine the Cohen-Macaulay property of binomial edge ideals of trees. We begin by introducing the necessary notations.
Notation 5.1.
Let $G=(G,x_{0})$ be a rooted tree. We label vertices of $G$ as follows: Label the root vertex $x_{0}$ as $1$ and the vertex at level $1$ of $G$ as $2$ (since $G$ has only one vertex at level $1$ as $x_{0}$ is a pendant vertex). Label the vertices at level $2$ of $G$ as $23,\ldots,m_{2}+$ 2 from left to right, where $m_{2}$ is number of vertices at level $2$. Label the vertices at level $3$ of $G$ as $m_{2}+3,\ldots,m_{2}+2+m_{3}$ from left to right, where $m_{3}$ is number of vertices at level $3$, and so on.
Notation 5.2.
Let $\mu$ be the Lexicographic term order on $S$ induced by $x_{1}>\cdots>x_{n}>y_{1}\cdots>y_{n}$. Let $\sigma$ be term order on $S[t]$ given by for any two monomials $M_{1},M_{2}\in S$ and two integers $i,j\geq 0$, we have $M_{1}t^{i}<_{\sigma}M_{2}t^{j}$ if and only if $i<j$ or $i=j$ and $M_{1}<_{\mu}M_{2}$.
Remark 5.1.
Let $G$ be a tree. Given an admissible path $\pi:i=i_{0},i_{1},\ldots,i_{n}=j$, with $i<j$ one has $i_{k}<j$ for all $0\leq k<n$. With respect to the above labeling on a tree $G$, one has every monomial $M$ in minimal generating set of $\operatorname{in}_{\mu}(J_{G})$ is divisible by some $x_{i}$ and if $x_{i}\mid M$ then $x_{j}\nmid M$ for all $j\neq i$ (cf. [10, Theorem 2.1]).
Notation 5.3.
Let $\Delta$ be a simplicial complex such that the facet ideal $I(\Delta)=\operatorname{in}_{\mu}(J_{G})$.
Definition 5.1.
[10, Section 10.3.4] Let $\Delta$ be a simplicial complex.
  1. (1)
    A cycle is an alternating sequence of distint vertices of facets $v_{1},F_{1},\ldots,F_{n},v_{n+1}=v_{1}$ such that $v_{i},v_{i+1}\in F_{i}$ for $i=1,\ldots,n$.
  2. (2)
    A cycle is special if it has no facet containing more than two vertices of the cycle.
Theorem 5.1.
Let $G$ be a tree and $J_{G}$ be the binomial edge ideal. Let $\Delta$ be a simplicial complex such that the facet ideal $I(\Delta)=\operatorname{in}_{\mu}(J_{G})$. Then one has
  1. (a)
    $\Delta$ does not have any special odd cycles.
  2. (b)
    $I(\Delta)$ is normally torsionfree.
  3. (c)
    $\operatorname{in}_{\mu}J_{G}^{(k)}=\operatorname{in}_{\mu}J_{G}^{k}$ for all $k\geq 1$.
  4. (d)
    $J_{G}^{(k)}=J_{G}^{k}$ for all $k\geq 1$.
  5. (e)
    $\mathcal{R}(\operatorname{in}_{\mu}J_{G})$ is Cohen Macaulay.
  6. (f)
    $\mathcal{R}(J_{G})$ is Cohen Macaulay.
Proof.
First, we relabel the vertices of $G$ as in Notation 5.1. Then, set $\{i_{1},\ldots,i_{n},j_{1},\ldots,j_{n}\}$ be the vertices of $\Delta$ such the variables $x_{k}$ and $y_{k}$ correspond to $i_{k}$ and $j_{k}$ in the associated Stanley-Reisner ring $S/I(\Delta)$, respectively. Observer that every facet of $\Delta$ has only one $i_{k}$’s. Each admissible path of length gearter than $1$ in $G$ is either between two distinct vertices of same level or between vertex $u$ at level $i$ and a vertex $v$ at level $i+1$ such that vertices in $N_{G}(v)$ at level $i$ are less than $u$. Therefore, $\Delta$ does not have any odd cycle which have facets containing atmost two vertices of the cycle, as desired. Since $\Delta$ does not have any special odd cycles, it follows from [9, Theorem 10.3.16] that $I(\Delta)$, that is, $\operatorname{in}_{\mu}(J_{G})$, is normally torsionfree. Then the assertions, (c) follows from [9, Theorem 10.3.15], (d) follows from [6, Theorem 3.3] and (e) follows from [9, Theorem 10.3.13]. Then from [6, Equation (2)] and the proof of [6, Lemma 3.1] it follows that $\operatorname{in}_{\mu}(J_{G})^{k}=\operatorname{in}_{\mu}(J_{G}^{k})$ for all $k\geq 1$. Therefore from [4, Theorem 2.7] it follows that $\operatorname{in}_{\sigma}(\mathcal{R}(J_{G}))=\mathcal{R}(\operatorname{in}_{\mu}J_{G})$. Therefore, from (e) it follows that $\operatorname{in}_{\sigma}(\mathcal{R}(J_{G}))$ and $\mathcal{R}(J_{G})$ are Cohen Macaulay. ∎
Remark 5.2.
In general, initial ideals of binomial edge ideals of unicyclic graphs are not normally torsionfree. For instance, $\operatorname{in}_{\mu}(J_{C_{5}})$ is not a normally torsionfree (cf. [6, Section 1]).
@@ 6. Counterexample
Acknowledgement. The first author thanks A. V. Jayanthan for his valuable comments. The second author gratefully acknowledges partial funding received from the Science and Engineering Research Board (SERB), India, under the Core Research Grant (CRG/2023/007668).

### v2

@@ Abstract.
An ideal $I$ of a commutative ring $R$ is said to be of linear type when its Rees algebra and symmetric algebra exhibit isomorphism. In this paper, we investigate the conjecture put forth by Jayanthan, Kumar, and Sarkar in [11] that if $G$ is a tree or a unicyclic graph, then the binomial edge ideal of $G$ is of linear type. Our investigation validates this conjecture for trees. However, our study reveals that not all unicyclic graphs adhere to this conjecture.
@@ Introduction
is called binomial edge ideal of the graph $G$. We call the binomial $f_{ij}$ associated to an edge $\{i,j\}\in E(G)$ as an edge binomial. In [7], Herzog et al. observed that the binomial edge ideal appears naturally in the study of conditional independence ideals, providing a valuable tool to investigate robustness theory in the context of algebraic statistics. In recent years, researchers have investigated various algebraic properties, including radicality, primality, Gröbner bases, primary decomposition, Cohen-Macaulayness, symbolic powers, among others, of binomial edge ideals in relation to their underlying combinatorial properties; see for partial list [5, 7, 14, 15, 18] and the references therein.
@@
[11, Conjecture 4.17. (a)] If $G$ is a tree or a unicyclic graph, then $J_{G}$ is of linear type.
@@
This paper is organized as follows. The following section recalls definitions, notations, and known results required in the rest of the paper. In Section 2, we introduce and study $p$-sequence ideals. In Section 3, we shown that binomial edge ideals of trees and class of unicyclic graphs form a $p$-sequence. We study the Rees algebra of binomial edge ideals of trees in Sections 4. In Section 5, we present a counterexample to Conjecture A for a class of unicyclic graphs.
@@ 1. Preliminaries
  • [14, Definition 3.1] Let $e^{\prime}=\{i,j\}\notin E(G)$ be an edge in $G\cup\{e^{\prime}\}$. Then $G_{(e^{\prime})}$ is the graph on vertex set $V(G)$ and edge set
@@
[11, Lemma 4.1] [7, Corollary 2.2] Let $G$ be a simple graph and $f$ be an element of $S$. Then $J_{G}:f=J_{G}:f^{n}$ for all $n\geq 2$.
@@
[14, Theorem 3.4] Let $G$ be a simple graph. Let $e=\{i,j\}\notin E(G)$ be a bridge in $G\cup\{e\}$. Then $J_{G}:f_{e}=J_{G_{e}}$.
@@
[14, Theorem 3.7] Let $G$ be a simple graph and $e=\{i,j\}\notin E(G)$ be an edge. Then
@@
[18, Corollary 2.5] For any bipartite graph $G$, one has $\beta_{i,i+1}(S/J_{G})=0$, for all $i$.
@@
[10, Definition 1.1] Let $A$ be a commutative ring. Set $z_{0}=0$. A sequence of elements $z_{1},\ldots,z_{m}$ in $A$ is said to be a $d$-sequence if it satisfies the conditions:
@@
[11, Proposition 4.10] Let $A$ be a polynomial ring over a field $\mathbb{K}$. Let $I\subset A$ be a homogeneous almost complete intersection ideal. Then $I$ is generated by a homogeneous $d$-sequence $z_{1},\ldots,z_{h+1}$ such that $z_{1},\ldots,z_{h}$ is a regular sequence, $h=\operatorname{ht}(I)$.
@@
[17, Definition 1.1] A relation on an ordered sequence of elements $z_{1},\ldots,z_{n}$ of a commutative ring $A$ is a form of positive degree in the graded polynomial ring $R=A[X_{i}\mid 1\leq i\leq n]$, where elements of $A$ have degree 00 and each $X_{i}$ has degree $1$, that yields 00 when evaluated at $z_{1},\ldots,z_{n}$. Let $\Lambda_{j}$ denote the ideal of $R$ generated by the relations on $z_{1},\ldots,z_{n}$ of degree at most $j$. Let $r$ be the least positive integer such that $\Lambda_{r}=\cup_{j\geq i}\Lambda_{j}$. The invariant $r$ is called the relation type of the ideal $I=(z_{1},\ldots,z_{n})$. An ideal of relation type $1$ is said to be of linear type.
@@
[4, p. 257] Let $A$ be a commutative ring. An ordered sequence of elements $z_{1},\ldots,z_{n}$ of a ring $A$ is said to be a sequence of linear type if each of the ideal $(z_{1},\ldots,z_{i})$ is of linear type for all $1\leq i\leq n$.
@@
    Consider a sequence of elements $\{x_{1}x_{3}x_{4}x_{5},x_{1}^{2}x_{2}x_{6},x_{1}^{2}x_{2}^{2}x_{3}x_{5}\}$ in $\mathbb{K}[x_{1},\ldots,x_{6}]$. In the specified order, this sequence readily forms a $d$-sequence. However, no permutation of this sequence constitutes a $p$-sequence. This assertion is verified by Computation 5.1.
@@
Suppose the sequence of elements $z_{1},\ldots,z_{n}$ in $S$ form a $d$-sequence. Then, it satisfies Equation (2.3), see [9, Proposition 3.1]. However, this property does not hold for $p$-sequences. For instance, consider a graph $G$ with edge set
@@ 4.1. Defining ideal of Rees algebra
The subsequent result follows from Theorem 3.1 and [11, Theorem 3.2], where we obtain a minimal generating set for the defining ideal of $\mathcal{R}(J_{G})$ when $G$ is a tree.
@@ 5. Counterexample
Acknowledgement. The first author thanks A. V. Jayanthan for his valuable comments and for pointing out an error in an earlier version. The second author gratefully acknowledges partial funding received from the Science and Engineering Research Board (SERB), India, under the Core Research Grant (CRG/2023/007668).